Intro 8 years ago, a new machine cost $8 million to purchase and an additional $490,000 for the installation. The machine wasPart 3 18 - Attempt 1/5 for 10 pts. What is the net salvage value? No decimals Submit
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

B Machine cost Intallation cost Total Cost Useful life $8,000,000.00 $490,000.00 $8,490,000.00 Annual depreciation $566,000 C

Cell reference -

AB Machine cost Intallation cost Total Cost Useful life 8000000 490000 =SUM(C2:03) 15 =C4/C5 =C4-C7*8 7 1 Annual depreciation

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.